Suppose an increase in price decreases quantity demanded from 210 to 190. Using the mid-point formula, the percentage change in quantity demanded is:
Answer #1

Answer:

% change in quantity demanded

= ( Q2 - Q1 ) / (Q2+Q1) / 2

= ( 190 - 210 ) / (190 + 210 ) / 2

= - 20 / 200

= - 0.1

= 10% decrease in quantity demanded